# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ 6 Conclusion
The innovation of labor-saving technologies occurred with the renewal of traditional extraction methodology in the coal mining sector in interwar Japan. These technology shocks induced the revision of labor regulations for female miners, which led to institutional change concerning the females’ local labor market in the coal mining area. In this study, I examine the impact of coal mines on the regional economy during historical industrialization in interwar Japan and investigate how the institutional change influenced the effect.
In order to avoid selection bias, I utilize data on demographics, early-life health, and labor outcomes constructed using a set of official census-based statistics. I exploit the exogenous variation from the geological stratum to structurally identify the impacts of the coal mines on the regional economy. Several robustness checks confirm that my main results are not sensitive to the potential confounding factors and influential observations and the use of alternative variation in the estimation.
I find that the coal mines increased the local population, which was greater in the sample after the full enforcement of the revised labor regulations. Evidence from the analyses using the local demographics indicates that the regulations may have forced female miners to exit the labor market and form their families. Before the full enforcement of the regulation, the occupational hazards for female miners substantially increased the early-life mortality via the biological mortality selection mechanism in utero. This institutional change, however, further led to improvements in early-life mortality.
The literature on institutional interventions in the labor market has provided evidence that labor regulations can impede regional developments as they can reduce the local labor supply and increase unemployment (Botero et al. 2004). The revised labor regulations reduced the female labor supply and led to the gender-biased structural shift in the case of interwar Japan. My main findings, however, show that technology-induced labor regulations accelerated the local population growth through the family formation channel. In this process, the regulations played a role in mitigating the occupational hazards for female miners and improving the early-life health status in the local economy. I show the role of institutional change from the gender-specific labor regulations in understanding the mechanisms behind regional developments.
@@
A.1 Female Miners in Coal Mine
By the end of the 19th century, many mines in the Chikuhō coalfields had finished mechanizing the processes involved in hauling, such as drainage pumps and hoisting machines. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, such mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the face (Sumiya 1968, pp.  310–312).
Figure A.1a shows a photo of miners at the mine face (kiriha) in the 1910s. A male skilled miner (saki-yama) extracted coal, whereas a female miner (ato-yama) brought the coal using a bamboo basket called sura. Figure A.1b shows that ato-yama tied a rope around her body to prevent the heavy sura from slipping off in the steep pits. The environment in the pit was generally poor. The air was polluted with dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. [^48] Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926 (Japan Mining Association 1928, p. table.6). Moreover, accidents caused by falling rocks occurred frequently. [^49] Available statistics indicate that the average number of accidents per year that occurred in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. (Japan Mining Association, 1928, p. table.6)..
@@ A.1 Female Miners in Coal Mine
The introduction of new mining technologies, such as coal cutters, was gradually adopted in the 1920s. The introduction of these technologies was accompanied by a transition from the traditional room-and-pillar mining method to the longwall method, which is more suitable for mechanical coal extraction. [^50] Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of larger machinery (Nishinarita 1985, p. 91). The number of miners, both male and female, declined during this transition. However, this did not entirely change the labor patterns of female miners. Approximately one in five miners was a woman in 1930 (Table 1), and they still worked inside the pits as ato-yama.
A significant event that changed the labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after World War I, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night, but the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. Table 1 shows that the number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were enacted in September. Mining companies responded to this decline by introducing coal conveyors (Tanaka and Oghino 1977, p. 79). [^51] See also Tanaka (1984, chapter. 4) for a comprehensive summary of the relationships between mechanization process and changes in labor patterns in the pits. Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. [^52] Since female miners carried out the coal selection process, females were still in danger of being exposed to the toxicities in the wastewater, even outside the pits. More than 15,000 women still worked in the mines in the late 1930s, which represented approximately 10% of all the miners measured.
A.2 Mechanization and Female Miners
@@
The history illustrated in Online Appendix A.1 suggests that the female miners were more likely to be replaced via mechanization. If this anecdote captures the true story, the mines adopting many machines were less likely to rely on the female miners. To assess this point, I assess the relationship between the degree of mechanization and the use of female miners in coal mines. The systematic data capturing the mechanization in coal mines is unavailable. Fortunately, however, the Fukuoka Prefectural Statistics document the horsepower and the number of miners in the coal mines in Fukuoka, including the Chikuhō coalfield. Figure A.2 illustrates the raw relationship between the share of female miners and log-transformed horsepower in the 82 coal mines in Fukuoka in 1930. Overall, this confirms the negative correlation between both variables. The estimate from Tobit model dealing with several censored observations in Figure A.2 is $-4.197$, which is statistically significant at the $1$% level. This suggests that a one standard deviation increase in the log-transformed horsepower ($=2.0$) decreases the use of female mines by approximately $8.4$%, which accounts for more than half of the standard deviation of the share of female miners ($=15.6$). Thus, this result supports the idea that mechanization via the installation of new technologies has worked to replace female miners in the coal mines.
@@
Note: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
@@
A.3 Influence of Great Depression in Japan
The impact of the Great Depression was not substantial in Japan compared to that in other countries (Miyamoto 2008, pp.  56–57; Abe 2008, p. 106; Blumenthal 1981, pp. 46). Blumenthal (1981, p.43) shows that the index of industrial production in Japan did not decrease dramatically between 1929 (=100) and 1930 (=97). In contrast, the US and the UK suffered much more significant reductions (81 in the US and 92 in the UK) in the same period. Blumenthal (1981, p. 45) also reports that the unemployment rates between 1930 and 1935 in the US, the UK, and Japan were approximately 27%, 15%, and 6%, respectively. Consequently, Japan’s GNP growth rate (from 1929 to 1930) was positive at $1.1$%, whereas it was $-7.7$ and $-0.1$, respectively, in the US and UK. This evidence supports the validity of using the systematic data from the census and vital statistics of the 1930s. In the mining sector, while several mines in Chikuhō coalfield were documented to have ceased operations during the depression, these were temporary (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973). Figure 1 confirms that the reduction in coal production was marginal in the case of Japan. Overall, the depression mainly hit prices rather than production. While the price declines were relatively large in textiles, metals, and construction materials, however, the shocks influenced food and fuel prices less (Nakamura and Odaka 1989, p. 53).
@@ A.4 Shares of Coal and GSC Mines
Figure A.3a shows the number of coal and oil mines documented in the AHCs in 1931 and 1936. Similarly, Figure A.3b shows the numbers of GSC and other metal mines. As shown, approximately 90 (85)of mines extracting fuels (metals) were coal (GSC) mines. Figure A.3 shows the number of mines, not municipalities. Since several mines coexist in a municipality, the number of municipalities with mines reported in the main text is smaller than the number of mines.
@@
Males and females derive utilities from consumption ($c$) and having children ($b$). For a spouse $g\in\{m,f\}$, additively separable utilitu function is specified as follows:
@@
$\footnotesize{\begin{split}c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f},\\ h_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}.\end{split}}$ (11)
@@
Therefore, $\frac{\partial b}{\partial w_{f}}$ becomes negative under following condition:
@@
I excluded several individuals whose statistics were unavailable or unreliable. Three individuals whose length of service was longer than the number of years they had used the nursery are excluded. Six individuals had more days in care than days in operation. These individuals are also excluded because of suspected listing and transcription errors. Finally, we excluded three individuals for whom statistics on the number of days of childcare were not available. As a result, this table uses statistics on the 88 individuals among $100$ children originally listed. Sources: Noyori 2010, pp. 80–82. The average number of days closed is obtained from the Statistics Bureau of the Cabinet 1926, p. 18.
To what extent were childcare services available to female miners at that time? As described in Section 2, female miners could use the nurseries attached to some coal mines. Fortunately, Noyori (2010) provides a comprehensive study on the nurseries in coal mines using a unique archive of the nursery survey in a few mines in Chikuhō. According to the nursery’s regulations, employees can use the nursery during working hours (Noyori 2010, p. 78). However, evidence shows that children over 5 years old were unlikely to use the nurseries in mines (Noyori 2010, pp. 76–77). Moreover, the number of days using the nursery was quite small. For example, Table B.1 summarizes the utilization of the nursery based on the survey report on 88 children at the Yamauchi Coal Mine in 1915. Column (a) shows that mothers’ average working years are 2.4, whereas their average years of nursery use are 1.3. Column (c) indicates that the average number of annual days used was roughly 85 days (median is $57.5$ days), whereas the average number of annual operating days available from the mining census is 320 (column d). The years spent in daycare ($p_{c}$) can be considered the number of years the mother worked after the child was born. Thus, one can calculate the utilization rate for each child ($c$) as follows:
@@
where $n_{c}$ is the number of days using nersery. Column (e) indicates that this rate is estimated to be $20$%. Note that several mothers used the nursery for more than 200 days, meaning that the nursery must have been opened throughout the operating days. This rejects the possibility that the nursery could not be opened due to a lack of manpower. Despite this, one must be careful about the possibility that miners might not work fully during the operating days. Even if one assumes that miners worked only one-third of the operating days (i.e., $107$ days), however, the utilization rate is still $60$% .
In addition, childcare requires fees, which were usually deducted from miners’ paychecks (Noyori 2010, p. 109). This implies that some miners who could not use the nursery may use alternative childcare institutions. Another possible informal institution may be hiring babysitting labor (komori). Noyori (2010, pp. 47–48), however, shows that these babysitters were usually children from other households. The young babysitter had to carry the infant on her back up to the mine entrance. The mother had to climb up and down many times like a “mouse” from the pits to the entrance to feed her baby. This anecdotal evidence suggests that the quality of babysitting was not high enough to compensate for childrearing.
Overall, while the nursery was a useful institution for female miners, it only replaced a small part of the childcare work of female miners. Although babysitting might be available, it offers incomplete services, covering only a limited part of mothers’ childcare needs. This means that the marketization parameter $\gamma$ shall be considerably low in inequality 18. Therefore, optimal fertility should be negatively correlated with the relative wage ($w_{f}/w_{m}$). This predicts the relationship that the fertility rate would be lower before the female labor regulation but could be higher after the regulation because the relative wage declined substantially after the prohibition of female miners’ work in the pits (Section 2.2).
@@ C.1 Mine Deposit
Data on the location and type of mines are obtained from the official reports published by the Association for Harmonious Cooperation (1932, 1937). They document all mines with more than $50$ miners in October 1931 and October 1936. The documents list three types of mines: metal, coal, and oil. [^53] The metal mine includes minerals such as gold, silver, copper, blister copper, pig iron, steel, iron sulfide, lead, zinc, bismuth, arsenious acid, tin, mercury, chromium, manganese, and sulfide minerals. A smelting mine is also included in the metal mine category. Coal mines include a few mines mining lignite. Oil mines include oil, crude oil, gas, volatile oil, kerosene, light oil, machine oil, and heavy oil. In these documents, the mine’s location is indicated by the name of the municipality. Hence, the name of the municipality was used while geocoding to match the mine dataset with the datasets of the other variables. The shapefile used for geocoding was obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism (available at: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html (accessed on 23rd February 2021). Several municipalities were consolidated between the years 1930 and 1938. Thus, I used the municipality boundaries defined in 1938 and aggregate statistics for consolidated municipalities. A minimal number of municipalities were divided; these were excluded from the original dataset. To check all consolidations and divisions of municipalities, I used Zenkoku shichōsonmei hensen sōran (Municipal Autonomy Research Association 2006), the most comprehensive book on the history of municipalities’ names.
@@
The initial dataset included both coal and heavy metal (gold, silver, and copper) mines. [^54] This is because the most representative heavy metals at that time in the Japanese mining sector were gold, silver, and copper (Online Appendix D.9). Therefore, to prepare the analytical samples, I had to exclude the duplications to better identify the impacts of coal mines. Panel A of Table C.1 summarizes duplications in the coal mine sample. Columns (1) and (2) show the figures based on the 1931 (1936) AHC, which is used to match with the 1930 (1935) Population Census and 1933 (1938) Vital Statistics datasets. Initially, there are $11,151$ municipalities in columns (1) and (2); note that the municipality boundaries are defined based on the 1938 boundaries. The number of municipalities within 30km of a coal mine is $1,147$ and $1,386$ in columns (1) and (2), respectively. The number of duplications in the treatment groups based on the $5$ km radius definition is 00 and $2$ in columns (1) and (2), respectively. Similarly, the number of duplications in the control groups based on the $5$ km radius definition is $7$ and $20$ in columns (1) and (2), respectively. Following the trimming step shown in Figure C.1a, the analytical coal mine samples are $1,140$ ($=1,147-7$) and 1,364 ($=1,386-22$), respectively. Panel B of Table C.1 summarizes the duplications in the GSC mine sample. Following the same procedure as in Figure C.1b, the GSC mine samples used in the robustness analysis are $2,498$ ($=2,510-22$) and $3,293$ ($=3,337-44$), respectively.
@@
The red (blue) circles in Figure C.2 indicate the coal (GSC) mine locations from the AHC. Figure 2 illustrates the spatial distribution of the coal mine sample by year and group. Figure 3 shows the distribution in Chikuhō coalfield in Kyushū in detail.
@@
Notes: These figures show how to trim the municipalities into the analytical samples. Figures C.1a and C.1b show the trimming steps for the coal and GSC (gold, silver, and copper) mines, respectively.
@@
Notes: The red and blue circles indicate the locations of coal and GSC (gold, silver, and copper) mines, respectively. Chikuhō coalfield in Fukuoka prefecture in Kyushū region is indicated with an arrow in Figure C.2a. Figure 3 illustrates the coal mines around Chikuhō coalfield in detail.
Source: Created by the author.

### v14

@@ 6 Conclusion
The innovation of labor-saving technologies occurred with the renewal of traditional extraction methodology in the coal mining sector during interwar Japan. These technology shocks induced the revision of labor regulations, leading to institutional change concerning the female mines. In this study, the impact of coal mines on the regional economy during historical industrialization in interwar Japan is examined to investigate how the institutional change influenced the effect.
In order to avoid selection bias, a set of official census-based statistics was used. The exogenous variation from the geological stratum is exploited to structurally identify the impacts of the coal mines on the regional development. Several robustness checks confirm that the main results are not sensitive to the potential confounding factors, influential observations, and the use of alternative variation in the estimation.
It is found that the coal mines increased the local population, and the estimated magnitude is greater in the sample after the full enforcement of the revised labor regulations. Evidence from the analyses using the register-based vital statistics indicates that the regulations had indeed forced female miners to exit the labor market and form families. Before the full enforcement of the regulation, the occupational hazards for female miners substantially increased the early-life mortality via the biological mortality selection mechanism in utero. The institutional change, however, further led to improvements in early-life mortality.
The literature on institutional interventions in the labor market has provided evidence that labor regulations can impede regional developments as they can reduce the local labor supply and increase unemployment (Botero et al. 2004). The revised labor regulations reduced the female labor supply and led to the gender-biased structural shift during interwar Japan. The findings, however, show that technology-induced labor regulations accelerated the local population growth through the family formation channel. In this process, the regulations played a role in mitigating the occupational hazards for female miners and improving the early-life health status in the local economy. The results demonstrate the role of institutional change from gender-specific labor regulations in understanding the mechanisms behind regional developments.
@@
A.1 Influence of Great Depression in Japan
The impact of the Great Depression was not substantial in Japan compared to other countries (Miyamoto 2008, pp. 56–57; Abe 2008, p. 106; Blumenthal 1981, pp. 46). Blumenthal (1981, p.43) shows that the index of industrial production in Japan did not decrease dramatically between 1929 (=100) and 1930 (=97). In contrast, the US and the UK suffered much more significant reductions (81 in the US and 92 in the UK) during the same period. Blumenthal (1981, p. 45) also reports that the unemployment rates between 1930 and 1935 in the US, the UK, and Japan were approximately 27%, 15%, and 6%, respectively. Consequently, Japan’s GNP growth rate (from 1929 to 1930) was positive at $1.1$%, whereas it was $-7.7$ and $-0.1$, respectively, in the US and UK. This evidence supports the validity of using the systematic data from the census and vital statistics of the 1930s. In the mining sector, while several mines in Chikuhō coalfield were documented to have ceased operations during the depression, these were temporary (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973). Figure 1 confirms that the reduction in coal production was marginal in the case of Japan. Overall, the depression mainly hit prices rather than production. While the price declines were relatively large in textiles, metals, and construction materials, the shocks influenced food and fuel prices less (Nakamura and Odaka 1989, p. 53).
A.2 Female Miners in Coal Mine
By the end of the 19th century, many mines in the Chikuhō coalfields had finished mechanizing the coal haulage processes. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, this mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312).
Figure A.1a shows a photograph of miners at the mine face (kiriha) in the 1910s. A male skilled miner (saki-yama) extracted coal, whereas a female miner (ato-yama) brought the coal to the surface using a bamboo basket called sura. Figure A.1b shows that ato-yama tied a rope around her body to prevent the heavy sura from slipping off in the steep pits. The environment in the pit was generally poor. Moreover, accidents caused by falling rocks occurred frequently. [^47] Available statistics indicate that the average number of accidents per year that occurred in the pits between 1917 and 1926, causing injuries and deaths of miners, was 142,595, of which 61,112 were from falling rocks. (Japan Mining Association, 1928, p. table.6). Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred 311 times per year on average between 1917 and 1926 (Japan Mining Association 1928, p. table.6)..
@@ A.2 Female Miners in Coal Mine
The introduction of new mining technologies, such as coal cutters, was gradually adopted in the 1920s. The introduction of these technologies was accompanied by a transition from the traditional room-and-pillar mining method to the longwall method, which is more suitable for mechanical coal extraction. [^48] Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of larger machinery (Nishinarita 1985, p. 91). The number of miners, both male and female, declined during this transition. However, this did not entirely change the labor patterns of female miners. Approximately one in five miners was a woman in 1930 (Table 1), and they still worked inside the pits as ato-yama. [^49] See also Tanaka (1984, chapter. 4) for a comprehensive summary of the relationships between mechanization process and changes in labor patterns in the pits. In short, introducing coal conveyors was the primary driving force for replacing the in-pit works of the female mines. This provided a strong incentive for mining companies to exclude female miners from the pits (Tanaka and Ogino 1977, p. 79). Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work. The number of female miners in the pits (ato-yama) dropped from 1930 to 1933, when the Revised Miner’s Labor Assistance Regulations were fully enacted in September (Table 1). Importantly, however, the revised regulations did not eliminate female miners altogether. Female miners were allowed to work in the pits of the thin layer and were primarily responsible for coal selection and other out-of-pit labor. More than 15,000 women still worked in the mines in the late 1930s, representing approximately 10% of all the miners measured.
A.3 Mechanization and Female Miners
@@
The history illustrated in Online Appendix A.2 suggests that female miners were more likely to be replaced via mechanization. If this anecdote is accurate, the mines deploying more machines were less likely to rely on the female miners. To assess this point, I examine the relationship between the degree of mechanization and the use of female miners in coal mines. The systematic data capturing mechanization in coal mines is unavailable. Fortunately, Fukuoka Prefectural Statistics document the horsepower and the number of miners in the coal mines in Fukuoka, including the Chikuhō coalfield. Figure A.2 illustrates the raw relationship between the share of female miners and log-transformed horsepower in the 82 coal mines in the prefecture in 1930. Overall, this confirms the negative correlation between both variables. The estimate from the Tobit model dealing with several censored observations in Figure A.2 is $-4.197$, which is statistically significant at the $1$% level. This suggests that a one standard deviation increase in the log-transformed horsepower ($=2.0$) decreases the use of female mines by approximately $8.4$%, which accounts for more than half of the standard deviation of the share of female miners ($=15.6$). Thus, this result supports the idea that mechanization via the installation of new technologies has worked to replace female miners in the coal mines.
@@
Note: This figure illustrates the raw relationship between the share of female miners and log-transformed horsepower in 82 coal mines in Fukuoka prefecture in 1930. The share of female miners is the number of female miners per 100 total miners. The mean and standard deviation of the share of female miners are $20.1$ and $15.6$, respectively. The horsepower is the total horsepower of the machines using any of steam, electrical, and water power. The mean and standard deviation of the log-transformed horsepower are $7.1$ and $2.0$, respectively. The red line represents the fitted values for a linear regression model. Although the Fukuoka Prefectural Statistics originally contained 90 coal mines, 8 mines were dropped because there were missing values in the horsepower statistics.
@@ A.4 Shares of Coal and GSC Mines
Figure A.3a shows the number of coal and oil mines documented in the AHCs in 1931 and 1936. Similarly, Figure A.3b shows the numbers of GSC and other metal mines. As shown, approximately 90 (85)of mines extracting fuels (metals) were coal (GSC) mines. Figure A.3 shows the number of mines. Since there were some cases where a few mines coexist in a municipality, the number of municipalities with mines reported in the main text is smaller than the number of mines.
@@
Males and females derive utilities from consumption ($c$) and having children ($b$). For a spouse $g\in\{m,f\}$, additively separable utility function is specified as follows:
@@
$\footnotesize{\begin{split}c_{m}+c_{f}+s=w_{m}n_{m}+w_{f}n_{f},\\ n_{g}+h_{g}=1~{}\text{for}~{}g\in\{m,f\}.\end{split}}$ (11)
@@
This is independent of the bargaining weight ($\theta$). Finally, equation 16 suggests that $\frac{\partial b}{\partial w_{f}}$ becomes negative under following condition:
@@
I excluded several individuals whose statistics were unavailable or unreliable. Three individuals whose length of service was longer than the number of years they had used the nursery are excluded. Six individuals had more days in care than days in operation. These individuals are also excluded because of suspected listing and transcription errors. Finally, I excluded three individuals for whom statistics on the number of days of childcare were not available. Consequently, this table uses statistics on the 88 individuals among $100$ individuals originally listed. Sources: Noyori 2010, pp. 80–82. The average number of days closed is obtained from the Statistics Bureau of the Cabinet 1926, p. 18.
To what extent were childcare services available to female miners at that time? As described in Section 2, female miners could use the nurseries attached to some coal mines. Fortunately, Noyori (2010) provides a comprehensive study on the nurseries in coal mines using a unique archive of the nursery survey in a few mines in Chikuhō coalfield. According to the nursery’s regulations, employees can use the nursery during working hours (Noyori 2010, p. 78). However, evidence shows that children over 5 years old were unlikely to use the nurseries in mines (Noyori 2010, pp. 76–77). Moreover, the number of days using the nursery was quite small. Table B.1 summarizes the utilization of the nursery based on the survey report on 88 children at the Yamauchi Coal Mine in 1915. Column (a) shows that mothers’ average working years are 2.4, whereas their average years of nursery use are 1.3. Column (c) indicates that the average number of annual days used was roughly 85 days (median is $57.5$ days), whereas the average number of annual operating days available from the mining census is 320 (column d). The years spent in daycare ($p_{c}$) can be considered the number of years the mother worked after the child was born. Thus, one can calculate the utilization rate for each child ($c$) as follows:
@@
where $n_{c}$ is the number of days using nursery. Column (e) indicates that this rate is estimated to be $20$%. Note that several mothers used the nursery for more than 200 days, meaning that the nursery must have been opened throughout the operating days. This rejects the possibility that the nursery could not be opened due to a lack of staff. Despite this, it is indeed possible that miners might not work fully during the operating days. Even if one assumes that miners worked only one-third of the operating days (i.e., $107$ days), the utilization rate is still $60$%.
In addition, childcare requires fees, which were usually deducted from miners’ paychecks (Noyori 2010, p. 109). This implies that some miners who could not use the nursery may use alternative childcare institutions. Another possible informal institution may be hiring babysitting labor (komori). Noyori (2010, pp. 47–48) shows that these babysitters were usually children from other households. The young babysitter was expected to carry the infant on her back up to the mine entrance. The mother would then climb up and down multiple times like a “mouse” from the pits to the entrance to feed her baby. This anecdotal evidence suggests that the quality of babysitting was not high enough to compensate for childrearing.
Overall, while the nursery was a useful institution for female miners, it only replaced a small part of the childcare work of female miners. Although babysitting was available, it offered an incomplete service, covering only a limited portion of mothers’ childcare needs. This meant that the marketization parameter $\gamma$ would be considerably low in inequality 18. Therefore, optimal fertility should be negatively correlated with the relative wage ($w_{f}/w_{m}$). This predicts the relationship that the fertility rate would be lower before the female labor regulation but could be higher after the regulation. This is because the relative wage declined substantially after the prohibition of female miners’ work in the pits (Section 2.2).
@@ C.1 Mine Deposit
Data on the location and type of mines are obtained from the official reports published by the Association for Harmonious Cooperation (1932, 1937). They document all mines with more than $50$ miners in October 1931 and October 1936. The documents list three types of mines: metal, coal, and oil. [^50] The metal mine includes minerals such as gold, silver, copper, blister copper, pig iron, steel, iron sulfide, lead, zinc, bismuth, arsenious acid, tin, mercury, chromium, manganese, and sulfide minerals. A smelting mine is also included in the metal mine category. Coal mines include a few mines mining lignite. Oil mines include oil, crude oil, gas, volatile oil, kerosene, light oil, machine oil, and heavy oil. In these documents, the mine’s location is indicated by the name of the municipality. Hence, the name of the municipality was used while geocoding to match the mine dataset with the datasets of the other variables. The shapefile used for geocoding was obtained from the official database of the Ministry of Land, Infrastructure, Transport and Tourism (available at: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html (accessed on 23rd February 2021). Several municipalities were consolidated between the years 1930 and 1938. Thus, the municipality boundaries defined in 1938 were used and aggregate statistics for consolidated municipalities. A minimal number of municipalities were divided; these were excluded from the original dataset. To check all consolidations and divisions of municipalities, Zenkoku shichōsonmei hensen sōran (Municipal Autonomy Research Association 2006), was used because it is the most comprehensive book on the history of municipalities’ names.
@@
The initial dataset included both coal and heavy metal (gold, silver, and copper) mines. [^51] This is because the most representative heavy metals at that time in the Japanese mining sector were gold, silver, and copper (Online Appendix D.9). Therefore, to prepare the analytical samples, the duplications were excluded to better identify the impact of coal mines. Panel A of Table C.1 summarizes duplications in the coal mine sample. Columns (1) and (2) show the figures based on the 1931 (1936) AHC, which is used to match with the 1930 (1935) Population Census and 1933 (1938) Vital Statistics datasets. Initially, there are $11,151$ municipalities in columns (1) and (2); note that the municipality boundaries are defined based on the 1938 boundaries. The number of municipalities within 30km of a coal mine is $1,147$ and $1,386$ in columns (1) and (2), respectively. The number of duplications in the treatment groups based on the $5$ km radius definition is 00 and $2$ in columns (1) and (2), respectively. Similarly, the number of duplications in the control groups based on the $5$ km radius definition is $7$ and $20$ in columns (1) and (2), respectively. Following the trimming step shown in Figure C.1a, the analytical coal mine samples are $1,140$ ($=1,147-7$) and 1,364 ($=1,386-22$), respectively. Panel B of Table C.1 summarizes the duplications in the GSC mine sample. Following the same procedure as in Figure C.1b, the GSC mine samples used in the robustness analysis are $2,498$ ($=2,510-22$) and $3,293$ ($=3,337-44$), respectively.
@@
The red (blue) circles in Figure C.2 indicate the coal (GSC) mine locations from the AHC. Figure 2 illustrates the spatial distribution of the coal mine sample by year and group. Figure 3 shows the distribution in Chikuhō coalfield in Kyūshū in detail.
@@
Notes: These figures demonstrare the procedure to trim the municipalities into the analytical samples. Figures C.1a and C.1b show the trimming steps for the coal and GSC (gold, silver, and copper) mines, respectively.
@@
Notes: The red and blue circles indicate the locations of coal and GSC (gold, silver, and copper) mines, respectively.
Source: Created by the author. Data on the location of mines are from the Association for Harmonious Cooperation (1932, 1937).